OpenStudy (anonymous):

Given: 3(x – 4) = 33 Prove: x = 15

OpenStudy (anonymous):

Use the following problem to answer the question below: Given: 3(x – 4) = 33 Prove: x = 15 Statements Reasons 3(x – 4) = 33 Given 3x – 12 = 33 Distributive 3x – 12 + 12 = 33 + 12 Addition property of equality 3x + 0 = 45 C 3x = 45 D x = 15 Division Property of Equality